The measure of the smallest angle in a triangle is 50 degrees less than the measure of the largest angle. The measure of the remaining angle is 10 degrees mor than the measure of the smallest angle. What is the measure of each angle?
Answer by askmemath(368)

let the smallest angle = X
Middle angle = X+10
Largest Angle= X+50
Sum of all angles in a trianle = 180
So
X + X+10 + X+50 = 180
3X+60 = 180
3X = 180-60 = 120
3X = 120
X = 120/3 = 40degrees
Middle angle = 40+10 = 50
Largest angle = 40+50 = 90